OpenStudy (anonymous):

Simplify square root of 3 multiplied by the fifth root of 3 choices: 3 to the power of 1 over 10 3 to the power of 3 over 5 3 to the power of 9 over 10 3 to the power of 7 over 10

OpenStudy (anonymous):

\[\sqrt{x} = \sqrt[2]{x^1} =x^\frac{ 1 }{ 2 }\]

OpenStudy (anonymous):

\[x^a \times x^b = x ^{a+b}\]

OpenStudy (anonymous):

Two rules that will help you solve this.

OpenStudy (anonymous):

how do I do that though?

OpenStudy (anonymous):

Well, according to the rule, what is 3 squared?

OpenStudy (anonymous):

irrational I think

OpenStudy (anonymous):

It's the same as writing 3 to the power of a half

OpenStudy (anonymous):